You are the new hired Business Development Manager for Company XYZ. The company wants to launch a new service product for diabetic management in the Manila market. You have the total budget as of $100,000 for this operation. Average price on market for the same product is $42 per member per month (revenue).
Operational expense cost project to be at 55%.
Calculate for ROI for 12 months with break even point
1
Expert's answer
2018-06-05T08:51:08-0400
You have the total budget as of $100,000 for this operation. Average price on market for the same product is $42 per member per month (revenue). Operational expense cost project to be at 55%.
Break even point BEP = 100,000/(42*12*(1 - 0.55)) = 441 units.
ROI for 12 months is:
ROI = 42*12*(1 - 0.55)*441/100,000 = 1.